OpenStudy (anonymous):

Evaluate the function rule for the given value. y = 15*3^x , for x = -3

OpenStudy (anonymous):

A) 5/3 B) 5/27 C) 135 D) 5/9

OpenStudy (anonymous):

@Tazmaniadevil

OpenStudy (anonymous):

5/9

OpenStudy (mathstudent55):

y = 15*3^x , for x = -3 Replace x with -3. What do you get?

OpenStudy (anonymous):

You have \(15 \times 3^{-3}\). In order to change the sign of an exponent, take the reciprocal, i.e. \(3^{-3} = \frac{1}{3^3}\). So now you have\[15\times 3^{-3} = 15\times \frac{ 1 }{ 3^3 } = \frac{15}{3^3}\]Now, evaluate \(3^3\), and simplify the fraction.

OpenStudy (anonymous):

15/27 @ospreytriple

OpenStudy (anonymous):

and then we simplify

OpenStudy (anonymous):

divide them by 3

OpenStudy (anonymous):

and we get 5/9

OpenStudy (anonymous):

I have about 2 same questions, I'll try to solve them and can u tell me if im right

OpenStudy (anonymous):

You are correct. Good job.